# Collaborative Aquatic Positioning System Utilising Multi-beam Sonar and Depth Sensors

### v2

@@ IV-A Simulation architecture
The CAP-SD system underwent evaluation within a custom Unity [29] underwater simulation environment. Portions of the simulation, such as sensors and communication framework, utilized MARUS [30], while the hydrodynamics components for ASVs and AUVs were implemented using the Dynamic Water Physics [31] asset in Unity.
Figure 4: CAP-SD simulation communication architecture
The simulation stack, which incorporates a ROS backend, is depicted in Fig 4. On the ROS [32] side, a gRPC server endpoint is established, functioning as a ROS node responsible for the bidirectional distribution of data flow. On MARUS side (Unity), it obtains inputs for actuators, mission control variables, and requests for simulating communication (either optical or acoustic) within the simulated environment from ROS.
IV-B Simulation Experiment Scenario
@@
Given that the application scenario of CAP-SD will be confined to indoor environments, the simulation scenario is set within a typical nuclear fuel pool (NFP). The entire area measures 28 m in length, 16 m in width, and has a depth of 8 m, as depicted in Fig 3 A and B. Within the NFP, the MallARD [28] (shown in Fig 3C and D), an ASV developed by the University of Manchester and equipped with simulated multi-beam sonar, IMU, and LiDAR, is utilized alongside the commercially available BlueROV2 (shown in Fig 3E and F), which serves as the underwater robot that is to be positioned. This strategic selection is aimed at facilitating the transition of CAP-SD from simulation to real-world deployment. It also lays the foundation for advancing technologies associated with underwater digital twins.
Figure 5: Simulation environment scenario and robots. (A) Overview of the simulated nuclear fuel pool  [^3] The CAD model of nuclear fuel pool used in this simulation is derived from the underground section (Bomb site B https://www.youtube.com/watch?v=wzbfPDgN5lo) of a map from the online game Counter-Strike 2, Nuke.: https://steamcommunity.com/sharedfiles/filedetails/?id=2324376136. (B) Sideoview of the nuclear fuel pool (B) Autonomous surface robot - MallARD (D) MallARD’s 3D model in Unity (E) Commercial underwater robot - BlueROV2 (F) BlueROV2 3D model in Unity.
IV-C Simulation evaluation of CAP-SD
@@
The simulation platform is a personal computer with an Intel Core i7 12700H CPU (central processing unit) running at 3.972 GHz and with 32-GB RAM (random access memory). In the experiments, the BlueROV2 moved in the tank while maintaining a fixed pitch and roll. In the following research, MallARD will be programmed to follow the BlueROV2 autonomously but the focus of the present research is to validate and quantify the accuracy of the proposed positioning system. All sensor data was collected via ROS and was recorded on the same roscore, which allows data to be shared in real-time and synchronised to a single clock.
@@
To evaluate the performance of the CAP-SD system, five distinct datasets with varying motion patterns were collected. In these five datasets, the BlueROV2 operates along five distinct trajectories: square, lawnmower, bouncing, random, and two-floor. All datasets were recorded when MallARD was held stationary and BlueROV2 was moved underneath in the multi-beam sonar’s FOV. Data from all sensors was recorded and was post processed using CAP-SD. Pose information from the Unity transformation component was recorded as the ground truth.
IV-D Discussion
@@
Figure 6: Plot showing the 3D trajectory of BlueROV2 estimated by CAP-SD against ground truth overtime respectively, during 71 seconds of dataset 5 (Two-floor).
@@
Figure 6B shows the 3D trajectory of the CAP-SD alongside the ground truth. Overall, within this simulation environment, CAP-SD is capable of accurately localising the ROV. Table 5 presents the RMS error of the proposed CAP-SD methods for all five datasets. Due to the depth/pressure plugin used in this simulation not actually sensing water pressure but merely adding noise to a known absolute depth, the results do not include a comparison of the RMSE along the $z$-axis. Therefore, as detailed in Section III, the accuracy of localisation largely depends on whether the YOLOv5 model can accurately determine the bounding box (the pixel coordinates of the robot on the sonar image frame).
As indicated by the Table 5, the Euclidean RMSE remains below 200 mm for each trajectory type. In datasets 2 and 3, the values are relatively higher, approaching 200 mm. This is attributed to the BlueROV2 passing through specific locations during the bouncing and lawnmower trajectories, where, in the sonar images, the spatial overlap between the robot and other objects in the underwater environment occurs. Under these circumstances, YOLOv5 fails to provide accurate pixel coordinates for the robot, as shown in Fig 7. In Fig 7A, the robot is not significantly affected by the environment, allowing YOLOv5 to clearly detect the BlueROV2 in the sonar images. However, in certain areas, as shown in Fig 7B, this is not the case. The green bounding box indicates the expected range, while the red represents the results from YOLOv5. This is also a challenge for multi-beam sonar: signal-to-noise ratio (SNR) and resolution. Images from multi-beam sonar frequently exhibit artifacts and possess a lower resolution when contrasted with images captured by cameras [33].
Figure 7: Sonar image view with YOLOv5 object detection. (A) Accurate robot detection by YOLOv5. (B) Due to YOLO v5’s inability to provide precise bounding boxes, errors occur in localisation. Green represents the true position, while red indicates the detection results of YOLOv5.
Figure 8: plots (A) and (C) shows the translation of the ROV in the the world fixed frame estimated using the two CAP-SD methods alongside the ground truth. Plots (B) and (D) show the associated error histograms for CAP-SD
@@
TABLE I: RMS errors relative to the ground truth measurement of CAP-SD for all 5 datasets  [^5] The source code and details about the rest dataset are available in: https://github.com/Xue1iang/IROS2024_CAP-SD.
Dataset Trajectory x RMSE y RMSE Euclidian RMSE
1
Square 134.2 mm 129.6 mm 160.1 mm
2
Bouncing 164.7 mm 144.5 mm 201.7 mm
3
Lawnmower 162.6 mm 147.3 mm 206.5 mm
4
Random 143.9 mm 134.7 mm 180.2 mm
5
Two-floor 167.7 mm 135.2 mm 197.4 mm
@@ V Conclusions and future works
This paper introduces a novel underwater localisation system for joint robotic operations, capable of sim-to-real applications. To the authors’ best knowledge, this is the first system applied in constrained underwater environments that does not require infrastructure, is not limited by lighting conditions, and is unaffected by water turbidity. Since it is in the proof-of-principle stage, stringent accuracy requirements have not been imposed.
This work has demonstrated the potential of the CAP system, however, there are still several challenges to overcome, including:
@@
  • A deeply customized neural network designed specifically to detect objects in sonar images
  • Real-world experimental evaluation
@@
  • Joint above-water and underwater 3D reconstruction in static/dynamic aquatic environments
  • Path planning for surface robots serving as external tracking systems to enhance the 3D SLAM capabilities of underwater robots
  • Temporal synchronisation of sensors distributed across multiple robotic agents
@@ Acknowledgment
This research was funded by EPSRC under grants: EP/P01366X/1, EP/W001128/1 and by an impact acceleration account secondment scheme which was jointly funded by The University of Manchester and EPSRC. Lennox acknowledges the support of the Royal Academy of Engineering (CiET1819/13).

### v3

@@ IV-A Simulation scenario
The CAP-SD system firstly underwent evaluation within a custom facility set up using the Unity [29] underwater simulation environment. Portions of the simulation, such as sensors and communication framework, utilized MARUS [30], while the hydrodynamics components for ASVs and AUVs were implemented using the Dynamic Water Physics [31] asset in Unity.
@@
Given that the application scenario of CAP-SD will be confined to indoor environments, the simulation scenario was set within a typical nuclear fuel pool (NFP). The entire area measured 28 m in length, 16 m in width, and had a depth of 8 m, as depicted in Fig 4 A and B. Within the NFP, the MallARD [28] (shown in Fig 4C and D), an ASV developed by the University of Manchester and equipped with simulated multi-beam sonar, IMU, and LiDAR, was utilized alongside the commercially available BlueROV2 (shown in Fig 4E and F), which served as the ROV that was to be positioned.
Figure 4: Simulation environment scenario and ROVs. (A) High-level view of a simulated nuclear fuel pool. (B) Autonomous surface robot - MallARD. (C) MallARD’s 3D model in Unity. (D) Commercial underwater ROV - BlueROV2 (E) BlueROV2 3D model in Unity.
IV-B Simulation evaluation of CAP-SD
@@
The simulation platform was a personal computer with an Intel Core i7 12700H CPU (central processing unit) running at 3.972 GHz and with 32-GB RAM (random access memory). In the experiments, the BlueROV2 moved in the tank while maintaining a fixed pitch and roll. In the following research, MallARD was programmed to follow the BlueROV2 autonomously, with the focus of the present research is to validate and quantify the accuracy of the proposed positioning system. All sensor data was collected via ROS and was recorded on the same roscore, which allowed data to be shared in real-time and synchronised to a single clock.
@@
To evaluate the performance of the CAP-SD system, five distinct datasets with varying motion patterns were collected. In these five datasets, the BlueROV2 manoeuvred itself along five distinct trajectories: square, lawnmower, bouncing, random, and two-floor. All datasets were recorded when MallARD was held stationary and BlueROV2 was moved underneath in the multi-beam sonar’s FOV. Data from all sensors was recorded and was post processed using CAP-SD. Pose information from the Unity transformation component was recorded as the ground truth.
IV-C Simulation results discussion
@@
Figure 5: The 3D trajectory of BlueROV2 estimated by CAP-SD against ground truth overtime respectively, during 71 seconds of dataset 5 (Two-floor).
Figure 6: The positioning simulation results of the CAP-SD system from the X, Y, and Z axes
@@
Figure 5B shows the 3D trajectory of the CAP-SD alongside the ground truth, for one of the motion patterns. Overall, within this simulation environment, CAP-SD was capable of localising the underwater ROV with reasonable accuracy. Table 3 presents the RMSE and mean euclidean distance (MED) of the proposed CAP-SD methods for all five datasets. For simplicity, the depth/pressure plugin that was used in this simulation, added noise to a known absolute depth, and hence the results do not include a comparison of the RMSE along the $z$-axis. Therefore, as detailed in Section III, the accuracy of localisation largely depends on whether the YOLOv5 model can accurately determine the bounding box (the pixel coordinates of the ROV on the sonar image frame). It is noted that when generating the real results, presented in Section XXX, the depth measurement was determined directly using a differential pressure cell attached to the ROV.
@@
TABLE I: RMSE and MED relative to the ground truth measurement of CAP-SD for all 5 datasets  [^3] All datasets are available in: https://github.com/Xue1iang/IROS2024_CAP-SD.
Dataset Trajectory x RMSE y RMSE MED
1 Square 134.2 mm 129.6 mm 160.1 mm
2 Bouncing 164.7 mm 144.5 mm 201.7 mm
3 Lawnmower 162.6 mm 147.3 mm 206.5 mm
4 Random 143.9 mm 134.7 mm 180.2 mm
5 Two-floor 218.3 mm 140.3 mm 228.3 mm
As indicated in Table 3, the MED remains below 200 mm for each trajectory type. In datasets 2 and 3, the values are relatively higher, approaching 200 mm. This is attributed to the BlueROV2 passing through specific locations during the bouncing and lawnmower trajectories, where, in the sonar images, the spatial overlap between the ROV and other objects in the underwater environment occurs. Under these circumstances, YOLOv5 fails to provide accurate pixel coordinates for the ROV, as shown in Fig 8. In Fig 8A, the ROV is not significantly affected by the environment, allowing YOLOv5 to clearly detect the BlueROV2 in the sonar images. However, in certain areas, as shown in Fig 8B and C, this is not the case. The green bounding box indicates the expected range, while the red represents the results from YOLOv5.
IV-D Experimental hardware architecture
In the experimental validation of CAP-SD, the ASV was MallARD. A multi-beam sonar (Oculus M3000d) was mounted onto the chassis of the ASV, while self-localisation relied on the Sick TiM571 LiDAR and Microstrain 3DMGQ7 IMU. The communication link to the base station was made tetherless. The underwater ROV used in thsi work was the tethered BlueROV2.
The CAP-SD system was assessed with data collected from a mock-up of a NFP (4.8 m $\times$ 3.6 m $\times$ 2.0 m), which was located at the RAICo1 lab in Whitehaven, UK. The positioning accuracy of the proposed technique was validated using the Qualisys motion tracking system, which consisted of six cameras, as depicted in Fig 7.
Figure 7: Hardware and System Architecture: Layout of MallARD components (left) and experimental field (right).
IV-E Online positioning performance
To evaluate the performance of CAP-SD when using real hardware, the ASV and ROV moved synchronously within the mock-up of the NFP. Under manual control, the BlueROV2 performed translational and rotational motions in the X-Y plane, with depth changes within a 2-meter range. Three sets of data, each lasting approximately 80 seconds, were collected during validation experiments. The differences among the data sets were minimal, mainly in recording times, with all experiments conducted under manual control. To better integrate sonar images as one of the inputs for real-time positioning in the CAP-SD system, the official Oculus software [32] was not used to display the sonar images (Fig 2B). Instead, the corresponding ROS package was adopted. Additionally, YOLOv5m was applied in real time on the ROV through the ROS-YOLO interface. Fig 9 presents a sample image of the BlueROV2 successfully detected by YOLO and an online visual representation of the ROV’s position.
IV-F Experimental results discussion
Fig 10A displays the online positioning results of CAP-SD compared to the ground truth. Due to significant overlap in the ROV’s trajectory over its entire dataset, it was not feasible to plot the full trajectory intuitively, so only 42 seconds of the trajectory are shown here (the full dataset can be found at [33]). Fig 10B(i), C(i), and D(i) show BlueROV2’s translation in the world frame, comparing CAP-SD and Qualisys (ground truth). Table II shows the full results for each dataset. The results indicate that the MED for CAP-SD ranged from 277.5 mm to 410.0 mm. Across all datasets, the MED was 331.1 mm. It was evident that the error along the $z$-axis was relatively small, which was due to the fact that depth information was primarily obtained from the depth sensor, which was highly accurate and the majority of the error resulted from estimation along the x and y axes. One of the reasons for this error was that in the CAP-SD system, YOLO detected and locked onto the centre of the object. However, when the BlueROV2 experienced changes in depth or planar movement, its roll and pitch angles varied. Consequently, the depth sensor, mounted at the tail of the BlueROV2, did not accurately reflect the depth of the ROV’s centre.
Figure 8: Sonar image view with YOLOv5 object detection. (A) Accurate ROV detection by YOLOv5 (B) and (C) Inaccurate ROV detection by YOLOv5.
Figure 9: Underwater experiment setup of CAP-SD and the corresponding FOV of the multi-beam sonar.
TABLE II: Comparative performance of CAP-SD across various datasets.
@@ V Conclusions and future works
This paper introduced a novel underwater localisation system for joint robotic operations, capable of sim-to-real applications. To the authors’ best knowledge, this is the first localisation system, applied in constrained underwater environments, that does not require infrastructure, is not limited by lighting conditions, and is unaffected by water turbidity. The paper validates the CAP-SD system through both simulation and experimental testing.
As well as demonstrating the potential of the CAP system, there are still several challenges that need to be overcome, including:
@@
  • A customized neural network designed specifically to detect objects in sonar images
@@
  • Using neural networks improves the CAP-SD system’s localisation capabilities, enabling it to determine the ROV’s 6-DOF pose.
  • Path planning for ASV serving as external tracking systems to enhance the underwater SLAM capabilities of ROV.
Figure 10: The positioning results of the CAP-SD system from the X, Y, and Z axes.A(i), B(i) and C(i) The positioning of BlueROV2 by CAP-SD in comparison with the ground truth along the X, Y, and Z axes, respectively. A(ii), B(ii) and C(ii) The error histograms of CAP-SD on the x, y, and z axes.
@@ Acknowledgment
This research was funded by EPSRC under grants: EP/P01366X/1, EP/W001128/1 and by an impact acceleration account secondment scheme which was jointly funded by the University of Manchester and EPSRC. Lennox acknowledges the support of the Royal Academy of Engineering (CiET1819/13). The authors would like to acknowledge the RAICo Fellowship Programme for providing support, resources, and valuable industry engagement opportunities, which significantly contributed to the development and application of this work.